# Efficient Generation of Multi-partite Entanglement between Non-local Superconducting Qubits using Classical Feedback

### v1

@@ V Entanglement Swapping and Teleportation
Similar to the teleportation-based CNOT protocol, long-distance state teleportation itself can also be implemented in constant depth. In Fig. 4, we outline a state-teleportation protocol from the perspective of quantum repeaters Jiang et al. (2009); Azuma et al. (2023), which serve as a helpful design to tackle imperfections during long-distance entanglement distribution in quantum networks Mastriani (2023). In a purely unitary circuit, swapping states between two ends of a register is performed by means of a cascade of SWAP gates (or alternatively, two unbounded fan-out gates; see Fig. 4a). However, by utilizing MCMs and feed-forward control, it is possible to instead teleport the state from one end of a register to another in constant depth by a quantum repeater protocol with simultaneous entanglement swapping Jiang et al. (2009). The structure of quantum repeaters is simple: ancillae qubits are prepared in a Bell state in an alternating manner, and Bell measurements are performed on overlapping qubits from each Bell state (see Fig. 4b). The measurement of each repeater can be executed in parallel, after which conditional feed-forward operations are applied to the qubit on the opposite side of the register from the qubit whose state we wish to teleport. Throughout this process, the ancillae qubits are involved in a process known as “entanglement swapping” Zukowski et al. (1993); Pan et al. (1998); Jennewein et al. (2001), since we get a long-distance entangled state from several short-distance ones. In fact, entanglement swapping generates a Bell pair distributed on two ends of the line of ancillae qubits (see Fig. 4c), and we can then use this Bell pair for teleportation task and other teleportation-based operations like CNOT we showed previously.
@@
Historically, entanglement swapping and teleportation were performed on photonic systems Schmid et al. (2009); Jin et al. (2015); Tsujimoto et al. (2018), NMR Nielsen et al. (1998), and atomic systems Barrett et al. (2004); Riebe et al. (2008). Deterministic teleportation has also been demonstrated on superconducting Steffen et al. (2013), but previous demonstrations of entanglement swapping on superconducting qubits Ning et al. (2019) were not deterministic. [^1] While Ref. Ning et al. (2019) claimed “deterministic” entanglement swapping, their protocol did not implement any feed-forward operations. Therefore, they were only able to deterministically prepare an entangled Bell state, without any active control over which state they prepared. Here, we demonstrate both deterministic state teleportation and deterministic remote entanglement generation via MCMs and feed-forward operations. In Fig. 4d, we plot the raw outputs for teleporting |$\ket{0}$, $\ket{+}$, and |$\ket{1}$ states from one side of our quantum processor to the other. The probability of measuring the incorrect state (calculated via the total variational distance to the ideal outcome) is 11.9%, 0.6%, and 5.6% for |$\ket{0}$, $\ket{+}$, and |$\ket{1}$, respectively. However, measuring the classical outputs of our teleportation protocol is not sufficient when the desire is to embed this sub-circuit within a larger unitary circuit, where maintaining coherence is important. For this purpose, in Fig. 4e we plot the Pauli transfer matrix (PTM) measured via quantum process tomography (QPT) for the state teleportation protocol. Ideally, the PTM should reflect the identity operation. While we measure a process fidelity of 0.67(1) — which is well above the classical threshold of 1/2 — the Pauli-$X$ and Pauli-$Y$ operators are not well-preserved by the process. In the PTM formalism, the preservation of a given Pauli eigenvalue is directly affected by non-commuting errors. Therefore, this is further evidence that our data qubits undergo strong measurement-induced dephasing during our MCMs, even with dynamical decoupling (see Appendix). Ultimately, this will limit the utility of the protocol within larger adaptive circuits.
@@
As previously stated, when performing state teleportation, we in fact entangle the two end ancillae qubits in a Bell state via entanglement swapping. We can therefore use this procedure as an alternative to the constant-depth GHZ state introduced in Sec. II, if entanglement is only desired between two non-local qubits. By varying the input state, [^2] The input to output mapping is: |$\ket{00}\mapsto\ket{\Phi^{+}}$, |$\ket{10}\mapsto\ket{\Phi^{-}}$, |$\ket{01}\mapsto\ket{\Psi^{+}}$, |$\ket{11}\mapsto\ket{\Psi^{-}}$. we can deterministically generate any Bell state between the two end ancillae. To demonstrate this, we deterministically prepare |$\ket{\Phi^{+}}=\tfrac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ and |$\ket{\Phi^{-}}=\tfrac{1}{\sqrt{2}}(\ket{00}-\ket{11})$, and once again measure their coherences via parity oscillations (shown in Fig. 4f). We measure Bell state fidelities of $F_{\ket{\Phi^{+}}}=0.57(1)$ and $F_{\ket{\Phi^{-}}}=0.55(1)$. While both of these are above the required threshold of 1/2 for genuine entanglement, it is clear from the contrast in the parity oscillations that a significant amount of phase coherence is lost in the process.
VI Outlook
@@
In this work, we have demonstrated a variety of protocols for generating non-local entangled states and entangling gates in constant depth. These protocols are based on teleportation-based quantum computing, whereby one has access to an entangled resource state which can be measured in the middle of a circuit, and whose classical outcomes can be rapidly analyzed in real-time to determine which conditional feed-forward operations to apply to the unmeasured qubits. For superconducting qubits, which have relatively fast gate times and relatively short coherence times compared to other hardware platforms, the utility of such protocols depends on the ability to perform MCMs and implement real-time feedback well within the coherence times of the qubits. With access to MCMs and fast classical feedback, teleportation-based protocols enable one to greatly reduce the circuit depth of many common tasks in quantum computing. This is particularly beneficial in the NISQ era, because it enables longer circuit depths to be performed within the coherence time of the qubits. Moreover, such protocols have similar requirements as quantum error correction (mid-circuit measurement, real-time syndrome decoding, etc.); therefore, developing the capabilities to implement teleportation-based protocols will be beneficial to the advancement of real-time quantum error correction.
@@
Nevertheless, the utility of such protocols for superconducting qubits is severely limited by the dephasing on idling spectator qubits during MCMs. This dephasing can result from the static $ZZ$ coupling between neighboring qubits Mitchell et al. (2021) during the measurement of one qubit, or it can result from any unintentional direct measurement of a qubit (e.g., due to frequency crowding of the readout resonators). Depending on the strength of the dephasing, methods such as dynamical decoupling may or may not help (see Appendix). In future work, we plan to explore ways to protect spectator qubits during MCMs. For example, in fixed-frequency systems, the static $ZZ$ coupling can be nulled by means of simultaneous off-resonant drives on both qubits Mitchell et al. (2021); or, in systems with tunable couplers, this term can be turned off entirely during measurement. To combat the unintentional (direct) measurement of a spectator qubit, we plan to explore the degree to which one can “cloak” the qubit from the readout cavity Lledó et al. (2023). Finally, because measurement- and teleportation-based protocols are fundamentally limited by the readout fidelity of MCMs, combining them with methods of performing readout correction of MCMs Hashim et al. (2023) would be an interesting avenue of exploration.
@@ Acknowledgements
A.H. and M.Y. devised the experiments and wrote the manuscript. A.H. performed the experiments and analyzed the data. M.Y. and P.G. provided theory support. L.C. and C.J. fabricated the sample. N.F., Y.X., and G.H. developed the classical control hardware used in this work. L.J. and I.S. supervised all work.
Competing Interests
All authors declare no competing interests.
@@ Appendix A Qubit & Readout Characterization
Figure A1: Quantum Processing Unit & Frequencies. (a) Eight-qubit superconducting transmon processor. Qubits are labeled in green, individual drive lines are labeled in blue, individual readout resonators (RO) are labeled in red, and the multiplexed readout bus (MRB) is labeled in cyan. The qubits are coupled to nearest neighbors in a ring geometry via coupling resonators (CR, purple). (b) Qutrit frequency spectrum. The solid lines and dashed lines denote the GE and EF transition of each qutrit, respectively. (b) Readout resonator frequency spectrum. The frequency of the readout resonator coupled to each qubit is shown by a dashed line.
The quantum processing unit (QPU) used in this work consists of eight superconducting transmon qubits arranged in a ring geometry (Fig. A1a). All of the qubits on the QPU can also be operated as qutrits. The GE and EF transition frequencies of each qutrit is plotted in Fig. A1b. Regions with frequency crowding can lead to microwave line crosstalk between transitions within the same subspace, or coherent leakage between transitions in different subspaces. For example, the EF transitions of Q0 and Q4 are close in frequency to the GE transition of GE; thus, driving Q2 and result in leakage on Q0 or Q4. A similar effect can occur between the GE transition of Q5 and the EF transition of Q6, which is spectrally far from the rest of the qubits on the QPU due to fabrication inaccuracies. The GE coherence times of each qubit is listed in Table A4.
Figure A2: Readout Calibration. Qutrit state discrimination is supported for all qubits on the quantum processor. MCM classification is determined by the in-phase ($I$) component of the readout signal. If $I>0$, the discriminator returns a 0; if $I<0$, the discriminator returns a 1. Therefore, MCMs cannot currently distinguish 1 from 2, and thus cannot capture leakage.
The readout calibration for all eight qutrits is plotted in Fig. A2, which displays the discrimination boundaries used to classify all results in this work. Qutrit state discrimination can be used to monitor leakage rates outside of the {|$\{\ket{0},\ket{1}\}$ computational basis; however, this is currently limited to final measurements, not MCMs (see Sec. C and Fig. A2). Alternatively, qutrit state discrimination can be used for improving qubit readout fidelities via excited state promotion (ESP) Mallet et al. (2009), whereby an $\pi_{1\rightarrow 2}$ pulse is applied to each qubit before readout, after which all |$\ket{2}$ state results are reclassified as |$\ket{1}$ in post-processing. ESP can protect qubits against $T_{1}$ decay or readout-induced decay during readout. We utilize ESP to improve qubit readout both in MCMs and final measurements. The readout fidelities with ESP are listed in Table A4.
@@ Appendix B Gate Benchmarking
The single-qubit and two-qubit gates used in this worked are benchmarked using randomized benchmarking (RB) and cycle benchmarking (CB) Erhard et al. (2019). Single-qubit Clifford infidelities measured via RB are listed in Table A4. Two-qubit Clifford and CZ gate infidelities — measured via RB and CB, respectively — are listed in Table A4. All quoted infidelities are given in terms of the process infidelity $e_{F}$, not the average gate infidelity $r$. These two are equal up to a simple dimensionality factor:
@@ Appendix C Classical Control Hardware
The QubiC gateware has a bank of distributed processor cores for performing pulse sequencing, parameterization, and conditional execution (i.e., control flow). The QubiC readout DSP (digital signal processing) chain includes on-FPGA demodulation and qubit state discrimination using a threshold mechanism. Currently, the discrimination is performed for MCMs using the in-phase ($I$) component of the integrated readout pulse. If $I>0$, the discriminator returns a 0; if $I<0$, the discriminator returns a 1. For this reason, all of the |$\ket{0}$ states are calibrated to be on the right side of $I=0$, and all of the |$\ket{1}$ and |$\ket{2}$ states are calibrated to be on the left side of $I=0$ (see Fig. A1c). These state-discriminated results can then be requested by any processor core (using a special instruction) and used as inputs to arbitrary control flow/branching decisions (e.g., a while loop or if/else code block). The total feedback latency (not including readout time) is 150 ns. In the future, more sophisticated discrimination will be available for MCMs capable of distinguishing |$\ket{1}$ from |$\ket{2}$.
Appendix D Measurement-Induced Dephasing

### v2

@@ V Entanglement Swapping and Teleportation
Similar to the teleportation-based CNOT protocol, long-distance state teleportation can also be implemented in constant depth. In Fig. 4, we outline a state-teleportation protocol from the perspective of quantum repeaters Jiang et al. (2009); Azuma et al. (2023), a useful design to tackle imperfections during long-distance entanglement distribution in quantum networks Mastriani (2023). In a purely unitary circuit, swapping states between two ends of a register is performed by means of a cascade of SWAP gates (or alternatively, two unbounded fan-out gates; see Fig. 4a). However, by utilizing MCMs and feed-forward control, it is possible to instead teleport the state from one end of a register to another in constant depth by a quantum repeater protocol with simultaneous entanglement swapping Jiang et al. (2009). The structure of quantum repeaters is simple: ancillae qubits are prepared in a Bell state in an alternating manner, and Bell measurements are performed on overlapping qubits from each Bell state (see Fig. 4b). The measurement of each repeater can be executed in parallel, after which conditional feed-forward operations are applied to the qubit on the opposite side of the register from the qubit whose state we wish to teleport. Throughout this process, the ancillae qubits are involved in a process known as “entanglement swapping” Zukowski et al. (1993); Pan et al. (1998); Jennewein et al. (2001), since we get a long-distance entangled state from several short-distance ones. In fact, entanglement swapping generates a Bell pair distributed on two ends of the line of ancillae qubits (see Fig. 4c), and we can then use this Bell pair for teleportation tasks and other teleportation-based operations like the CNOT we showed previously.
@@
Historically, entanglement swapping and teleportation were performed on photonic systems Schmid et al. (2009); Jin et al. (2015); Tsujimoto et al. (2018), NMR Nielsen et al. (1998), and atomic systems Barrett et al. (2004); Riebe et al. (2008). Deterministic teleportation has also been demonstrated on superconducting circuits Steffen et al. (2013), but previous demonstrations of entanglement swapping on superconducting qubits Ning et al. (2019) were not deterministic [^1] While Ref. 64 claimed “deterministic” entanglement swapping, their protocol did not implement any feed-forward operations. Therefore, they were only able to deterministically prepare an entangled Bell state, without any active control over which state they prepared.. Here, we demonstrate both deterministic state teleportation and deterministic remote entanglement generation via MCMs and feed-forward operations. In Fig. 4d, we plot the raw outputs for teleporting |$\ket{0}$, $\ket{+}$, and |$\ket{1}$ states from one side of our quantum processor to the other. The probability of measuring the incorrect state (calculated via the total variational distance to the ideal outcome) is 11.9%, 0.6%, and 5.6% for |$\ket{0}$, $\ket{+}$, and |$\ket{1}$, respectively. However, measuring the classical outputs of our teleportation protocol is not sufficient when the desire is to embed this sub-circuit within a larger unitary circuit, where maintaining coherence is important. For this purpose, in Fig. 4e we plot the Pauli transfer matrix (PTM) estimated via quantum process tomography (QPT) for the state teleportation protocol, which (in contrast to the previous protocols which had multiple data qubits and, thus, a larger chance to dephase) is feasible at the level of a single qubit. Ideally, the PTM should reflect the identity operation, where the input states are prepared on a qubit on one side of the register, and the output states are measured on a qubit on the other side of the register. While we measure a process fidelity of 0.67(1) — which is well above the classical threshold of 1/2 — the Pauli $X$ and $Y$ eigenvalues are not well-preserved by the process. In the PTM formalism, the preservation of a given Pauli eigenvalue is directly affected by non-commuting errors Hashim et al. (2024). Therefore, this is further evidence that our data qubits undergo strong measurement-induced dephasing during our MCMs, even with dynamical decoupling (see Appendix E). Ultimately, this will limit the utility of the protocol within larger adaptive circuits.
@@
As previously stated, when performing state teleportation, we in fact entangle the two end ancillae qubits in a Bell state via entanglement swapping. We can therefore use this procedure as an alternative to the constant-depth GHZ state introduced in Sec. II, if entanglement is only desired between two non-local qubits. By varying the input state [^2] The input to output mapping is: |$\ket{00}\mapsto\ket{\Phi^{+}}$, |$\ket{10}\mapsto\ket{\Phi^{-}}$, |$\ket{01}\mapsto\ket{\Psi^{+}}$, |$\ket{11}\mapsto\ket{\Psi^{-}}$., we can deterministically generate any Bell state between the two end ancillae. To demonstrate this, we deterministically prepare |$\ket{\Phi^{+}}=\tfrac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ and |$\ket{\Phi^{-}}=\tfrac{1}{\sqrt{2}}(\ket{00}-\ket{11})$, and once again measure their coherences via parity oscillations (shown in Fig. 4f). We measure Bell state fidelities of $F_{\ket{\Phi^{+}}}=0.57(1)$ and $F_{\ket{\Phi^{-}}}=0.55(1)$. While both of these are above the required threshold of 1/2 for genuine entanglement, it is clear from the contrast in the parity oscillations that a significant amount of phase coherence is lost in the process.
VI Conclusion and Outlook
@@
In this work, we have demonstrated a variety of protocols for generating non-local entangled states and entangling gates in constant depth. These protocols are based on teleportation- and measurement-based quantum computing, whereby one has access to an entangled resource state which can be measured in the middle of a circuit, and whose classical outcomes can be rapidly analyzed in real-time to determine which conditional feed-forward operations to apply to the unmeasured qubits. For superconducting qubits, which have relatively fast gate times and relatively short coherence times compared to other hardware platforms, the utility of such protocols depends on the ability to perform MCMs and implement real-time feedback well within the coherence times of the qubits. With access to MCMs and fast classical feedback, these protocols enable one to greatly reduce the circuit depth of many common tasks in quantum computing. This could be particularly beneficial in the NISQ era, because it enables longer circuit depths to be performed within the coherence time of the qubits. Moreover, such protocols have similar requirements as quantum error correction (mid-circuit measurement, real-time syndrome decoding, etc.); therefore, developing the capabilities to implement teleportation- and measurement-based protocols will be beneficial to the advancement of real-time quantum error correction.
@@
Nevertheless, the utility of such protocols for superconducting qubits is severely limited by the dephasing on idling spectator qubits during MCMs. This dephasing can result from the static $ZZ$ coupling between neighboring qubits Mitchell et al. (2021) during the measurement of one qubit Jurcevic and Govia (2022), or it can result from any unintentional direct measurement of a qubit (e.g., due to frequency crowding of the readout resonators). Depending on the strength of the dephasing, methods such as dynamical decoupling may or may not help (see Appendix E). In future work, we plan to explore ways to protect spectator qubits during MCMs. For example, in fixed-frequency systems, the static $ZZ$ coupling can be nulled by means of simultaneous off-resonant drives on both qubits Mitchell et al. (2021); or, in systems with tunable couplers, this term can be turned off entirely during measurement. To combat the unintentional (direct) measurement of a spectator qubit, we plan to explore the degree to which one can “cloak” the qubit from the readout cavity Lledó et al. (2023). Finally, because teleportation- and measurement-based protocols are fundamentally limited by the readout fidelity of MCMs, combining them with methods of performing readout correction of MCMs Hashim et al. (2023) would be an interesting avenue of exploration.
@@ Acknowledgements
Author Declarations
Conflict of Interest
All authors declare no conflict of interest.
@@
A.H. and M.Y. devised the experiments and wrote the manuscript. A.H. performed the experiments and analyzed the data. M.Y. and P.G. provided theory support. L.C. and C.J. fabricated the sample. N.F., Y.X., G.H., and K.N. developed the classical control hardware used in this work. L.J. and I.S. supervised all work.
@@
Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7
$T_{1}$ ($\mu$s) 96.6(2.6) 130.0(2.7) 142.0(3.0) 140.0(6.3) 77.0(5.2) 30.4(0.95) 55.6(1.3) 22.5(0.32)
$T_{2}^{*}$ ($\mu$s) 120.0(14.0) 41.0(7.2) 92.0(16.0) 61.0(6.1) 38.0(5.4) 8.5(1.3) 26.0(3.7) 39.0(1.7)
$T_{2E}$ ($\mu$s) 120.0(8.3) 130.0(7.5) 140.0(12.0) 90.0(13.0) 110.0(11.0) 33.0(3.6) 90.0(14.0) 43.0(2.2)
Table 1: Qubit Coherences. Qubit coherence times ($T_{1}$, $T_{2}^{*}$, $T_{2E}$) are listed above.
Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7
$P(0|0)$ 0.995(1) 0.995(1) 0.995(1) 0.992(2) 0.998(1) 0.991(1) 0.997(1) 0.987(3)
$P(1|1)$ 0.983(2) 0.962(7) 0.994(2) 0.986(2) 0.984(5) 0.986(2) 0.994(2) 0.986(2)
Table 2: Readout Fidelities. Simultaneous readout fidelities for all qubits with excited state promotion.
Q0 Q1 Q2 Q3 Q4 Q5 Q6 Q7
RB iso. ($10^{-3}$) 1.5(1) 0.33(2) 0.54(3) 1.0(1) 3.2(1) 1.92(9) 2.4(3) 2.58(9)
RB sim. ($10^{-3}$) 2.1(2) 3.1(2) 2.0(3) 1.9(2) 6.1(7) 5.7(3) 3.4(3) 7.6(9)
Table 3: Single-qubit Gate Infidelities. The process infidelities for isolated and simultaneous single-qubit gates measured via RB for each qubit are listed above.
(Q0, Q1) (Q1, Q2) (Q2, Q3) (Q3, Q4) (Q4, Q5) (Q5, Q6) (Q6, Q7) (Q7, Q0)
RB iso. ($10^{-2}$) 5.0(5) 1.62(7) 1.64(8) 2.6(2) 3.8(2) 4.6(2) 6.7(6) 3.8(3)
CB (CZ) ($10^{-2}$) 1.4(1) 0.57(1) 0.41(1) 0.81(4) 1.80(8) 2.08(3) 2.77(8) 1.40(7)
Table 4: Two-qubit Gate Infidelities. The process infidelities for two-qubit RB are listed above for each qubit pair used in this work. Native (CZ) gate fidelities are measured via CB.
@@ Appendix A Qubit & Readout Characterization
Figure 5: Quantum Processing Unit & Frequencies. (a) Eight-qubit superconducting transmon processor. Qubits are labeled in green, individual drive lines are labeled in blue, individual readout resonators (RO) are labeled in red, and the multiplexed readout bus (MRB) is labeled in cyan. The qubits are coupled to nearest neighbors in a ring geometry via coupling resonators (CR, purple). (b) Qutrit frequency spectrum. The solid lines and dashed lines denote the GE and EF transition of each qutrit, respectively. (b) Readout resonator frequency spectrum. The frequency of the readout resonator coupled to each qubit is shown by a dashed line.
The quantum processing unit (QPU) used in this work consists of eight superconducting transmon qubits arranged in a ring geometry (Fig. 5a). All of the qubits on the QPU can also be operated as qutrits. The |$\ket{0}\rightarrow\ket{1}$ (i.e., “GE”) and |$\ket{1}\rightarrow\ket{2}$ (i.e., “EF”) transition frequencies of each qutrit is plotted in Fig. 5b. Regions with frequency crowding can lead to microwave line crosstalk between transitions within the same subspace, or coherent leakage between transitions in different subspaces. For example, the EF transitions of Q0 and Q4 are close in frequency to the GE transition of Q2; thus, driving Q2 and result in leakage on Q0 or Q4. A similar effect can occur between the GE transition of Q5 and the EF transition of Q6, which is spectrally far from the rest of the qubits on the QPU due to fabrication inaccuracies. The GE coherence times of each qubit is listed in Table 4.
Figure 6: Readout Calibration. Qutrit state discrimination is supported for all qubits on the quantum processor. MCM classification is determined by the in-phase ($I$) component of the readout signal. If $I>0$, the discriminator returns a 0; if $I<0$, the discriminator returns a 1. Therefore, MCMs cannot currently distinguish 1 from 2, and thus cannot capture leakage.
The readout calibration for all eight qutrits is plotted in Fig. 6, which displays the discrimination boundaries used to classify all results in this work. Qutrit state discrimination can be used to monitor leakage rates outside of the {|$\{\ket{0},\ket{1}\}$ computational basis; however, this is currently limited to terminating measurements, not MCMs (see Sec. C and Fig. 6). Alternatively, qutrit state discrimination can be used for improving qubit readout fidelities via excited state promotion (ESP) Mallet et al. (2009), whereby an $\pi_{1\rightarrow 2}$ pulse is applied to each qubit before readout, after which all |$\ket{2}$ state results are reclassified as |$\ket{1}$ in post-processing. ESP can protect qubits against $T_{1}$ decay or readout-induced decay during readout. We utilize ESP to improve qubit readout both in MCMs and final measurements. The readout fidelities with ESP are listed in Table 4.
@@ Appendix B Gate Benchmarking
The single-qubit and two-qubit gates used in this worked are benchmarked using randomized benchmarking (RB) and cycle benchmarking (CB) Erhard et al. (2019). Single-qubit Clifford infidelities measured via RB are listed in Table 4. Two-qubit Clifford and CZ gate infidelities — measured via RB and CB, respectively — are listed in Table 4. All quoted infidelities are given in terms of the process infidelity $e_{F}$, not the average gate infidelity $r$. These two are equal up to a simple dimensionality factor:
@@ Appendix C Classical Control Hardware
The QubiC gateware has a bank of distributed processor cores for performing pulse sequencing, parameterization, and conditional execution (i.e., control flow) Fruitwala et al. (2024). The QubiC readout DSP (digital signal processing) chain includes on-FPGA demodulation and qubit state discrimination using a threshold mechanism. Currently, the discrimination is performed for MCMs using the in-phase ($I$) component of the integrated readout pulse. If $I>0$, the discriminator returns a 0; if $I<0$, the discriminator returns a 1. For this reason, all of the |$\ket{0}$ states are calibrated to be on the right side of $I=0$, and all of the |$\ket{1}$ and |$\ket{2}$ states are calibrated to be on the left side of $I=0$ (see Fig. 6). These state-discriminated results can then be requested by any processor core (using a special instruction) and used as inputs to arbitrary control flow/branching decisions (e.g., a while loop or if/else code block) Fruitwala et al. (2024). The total feedback latency (not including readout time) is 150 ns. After these experiments were performed, a neural network-based readout discriminator was developed for MCMs performed on QubiC which is capable of distinguishing |$\ket{1}$ from |$\ket{2}$ Vora et al. (2024).
Appendix D Look-up Tables
Outcome Conditional Operation
0
1 $X_{0}$ or $X_{2}$
Table 5: LUT for 2-qubit GHZ State.
Outcome Conditional Operation
00
01 $X_{4}$
10 $X_{0}$
11 $X_{2}$
Table 6: LUT for 3-qubit GHZ State.
Outcome Conditional Operation
000
001 $X_{6}$
010 $X_{0}X_{2}$ or $X_{4}X_{6}$
100 $X_{0}$
011 $X_{4}$
101 $X_{0}X_{6}$ or $X_{2}X_{4}$
110 $X_{2}$
111 $X_{0}X_{4}$ or $X_{2}X_{6}$
Table 7: LUT for 4-qubit GHZ State.
Outcome Conditional Operation
0, 0
0, 1 $Z_{0}$
1, 0 $X_{4}$
1, 1 $Z_{0}X_{4}$
Table 8: LUT for Teleportation-based CNOT. The bit-string results for the teleportation-based CNOT do not need to be analyzed holistically; rather, they are analyzed individually for each MCM (see Fig. 2a). Thus, we do not need to nest branch statements on our hardware.
Outcome Conditional Operation
0, 00
0, 01 $Z_{0}$
0, 10 $Z_{0}$
0, 11
1, 00 $X_{2}X_{4}$
1, 01 $Z_{0}X_{2}X_{4}$
1, 10 $Z_{0}X_{2}X_{4}$
1, 11 $X_{2}X_{4}$
Table 9: LUT for Fan-out Gate. The conditional $X$ gates are determined by the outcome of the first measurement, and the conditional $Z$ gate is determined by the combined outcome of the second two measurements.
In this work, look-up tables (LUTs) were utilized for decision branching based on the results of the MCMs. While LUTs are not scalable as system size increases, and will eventually need to be replaced by real-time decoders, they are still viable for small system sizes. The LUTs are constructed by nesting branch statements on QubiC using the distributed processor architecture Fruitwala et al. (2024). These branching statements poll the results of the measured ancillae qubits in a MCM, and then applies a conditional feed-forward operation on data qubit(s) depending on the results of the measurements.
For example, consider the task of preparing an $n$-qubit GHZ state using the protocol introduced in Sec. II. In the stabilizer formalism, a GHZ state is simultaneous eigenstate of $S_{0}=X_{1}X_{2}...X_{n}$ and $S_{j}=Z_{j}Z_{j+1}$ ($j=1,2,...,n-1$) with eigenvalue $+1$. Here, $S_{j}$ can be used to check the parity of any two neighboring qubits. Thus, the strategy is to initialize the data qubits in the $\ket{+}$ state so that it is an eigenstate of $S_{0}$ with an eigenvalue of $+1$, and then measure the stabilizers $S_{j}$ in parallel using a MCM on all of the ancillae qubits. The outcome of the stabilizer measurements are then analyzed holistically to determine on which data qubits we should perform a conditional bit-flip to reconstruct the $n$-qubit GHZ state. For example, suppose the data qubits are in the state |$\ket{\psi}=\tfrac{1}{2}\left(\ket{000110}+\ket{111001}\right)$ prior to measurement. By entangling the data qubits pairwise with ancillae qubits, as shown in Fig. 1, we can measure the stabilizers $\{S_{1},S_{2},S_{3},S_{4},S_{5}\}$. In the absence of any errors, a measurement of the stabilizers will return the values $\{S_{1},S_{2},S_{3},S_{4},S_{5}\}=\{0,0,1,0,1\}$, where a measurement of 00 ($1$) corresponds to a parity of $+1$ ($-1$); this tells us that we should apply a bit-flip ($X$ gate) on the 4th and 5th qubit to recover a GHZ state. Building up a LUT of all possible outcomes allows us to quickly apply conditional bit-flips in real-time, without the latency of having to decode the results on the FPGA or on some other (potentially non-local) classical processor. Though, it is worth noting that a real-time decoder of parity measurements could be implemented by analyzing the parity measurements sequentially (per clock cycle) and considering each 1 measurement as a domain wall which takes us from apply nothing to apply $X$, and back again.
In Tables 79, we include the LUTs used in this work for the various experiments performed in the main text. For each experiment, we list the various possible outcomes of the parity measurements, and for each outcome we list the possible conditional operations which will recover the intended target state (preference is given to the conditional operations with the minimal number of gates). Note that all data qubits are indexed from 0 in these tables. Note also that the bit-strings are read left-to-right in accordance with increasing ancilla qubit number label. We omit a LUT for the entanglement swapping and teleportation protocols (Sec. V), since, like the teleportation-based CNOT, we do nothing (apply the conditional gate operation) if we measure (0) 1.
Appendix E Measurement-Induced Dephasing